Determine the entropy change in the surroundings fhelp
Answer #1

dSsurroundings...

so..

S = Q/T since Q = -HRxn of system

so

Q = 906 kJ

T = 25°C = 298K

therefore

S surroundings = (906*10^3)/(298) =3040.268 J/molK

so..

S surroundings 3. 04 kJ/mnolK

choose B